# VITA: Zero-Shot Value Functions via Test-Time Adaptation of Vision–Language Models

### v4

@@ 4.2 Baselines
We compare VITA against value functions based on contrastive and autoregressive zero-shot VLMs. VLM-CL computes zero-shot value estimates from cosine similarity between CLIP representations of each frame and task description [baumli2023vlm]. VLM-RM regularizes CLIP embeddings by projecting features along the direction from a generic reference prompt to the task prompt [rocamonde2024vlm]. CLIP-FT trains a supervised regression head on frozen CLIP multimodal representations. CLIP-GRU trains a Gated Recurrent Unit (GRU) [cho2014gru] on frozen CLIP multimodal representations. GVL [ma2024vlm] is the state-of-the-art zero-shot value function that leverages autoregressive VLMs. In our experiments, GVL-0S refers to the zero-shot setting, where the VLM is prompted with only the test trajectory and task description. GVL-1S corresponds to a one-shot in-context setting, where a full shuffled trajectory from the training distribution, along with its progress labels, is provided as an example. We follow GVL in using Gemini 1.5 Pro (gemini-1.5 Pro-latest[team2024gemini] with its proposed prompt template. We also tested Qwen-VL 2.5 [qwen2.5], which failed to overcome temporal bias, and GPT-4o [hurst2024gptshort], which refused to produce value estimates. Additional implementation details are provided in Appendix D.
@@ 4.3 Evaluating Generalization Under Distribution Shifts
We evaluate the ability of progress estimators to generalize across novel environments, tasks, and robot embodiments using subsets of BridgeData V2 curated to introduce variation along these axes. Environment shifts involve changes in scene layout or background. For example, lm_pnp is a pick-and-place task in front of a laundry machine, while td_fold, ft_fold, and rd_fold feature cloth folding on different surfaces. ms_sweep introduces a long-horizon sweeping task in a confined tray. Embodiment shifts are evaluated using the DeepThought robot, which differs from the training embodiment (WidowX 250) in both morphology and camera perspective. dt_tk_pnp (pick-and-place) and dt_tk_stack (stacking) retain the in-distribution environment with a new embodiment, while dt_ft_stack (stacking) and dt_rd_pnp (drawer pick-and-place) involve both embodiment and environment shifts. Each dataset includes 200 expert trajectories with task descriptions, except for ms_sweep, which contains 100 due to its size limitation. Appendix B provides a full description of our evaluation datasets. We evaluate the performance of a value function estimator using the Value Order Correlation (VOC) metric [ma2024vlm], which measures the alignment between the predicted progress towards task completion and the chronological order of the frames in a visual trajectory. Formally, let $p_{1},\dots,p_{T}$ denote the predicted progress values for a trajectory of $T$ frames, and let $r_{t}=t$ represent the temporal index. VOC is defined as the Spearman rank correlation $\rho_{s}$ between the predicted values and the temporal indices.
@@
In Table 1, the results show that both VITA and CLIP-GRU consistently outperform GVL-0 and GVL-1, highlighting the limitations of pre-trained autoregressive VLMs and the importance of preserving temporal order for effective progress estimation. While both demonstrate effective progress estimation, VITA outperforms CLIP-GRU on the majority of evaluation datasets (6 out of 10). Both VITA and CLIP-GRU outperform CLIP-based baselines that lack temporal modeling (CLIP-FT, VLM-CL, VLM-RM), indicating the impact of encoding temporal history on progress estimation. CLIP-FT performs comparably to GVL on the in-distribution task but fails to generalize to out-of-distribution settings. Both VLM-CL and VLM-RM perform poorly, likely due to their lack of temporal modeling. For GVL-0S and GVL-1S, in some cases, in-context examples improve performance, while in others, they provide no benefit or even degrade performance. Both GVL methods perform well on folding tasks (td_fold, ft_fold, rd_fold), but fail to achieve comparable performance on stacking (dt_tk_stack, dt_ft_stack) and pick-and-place (dt_tk_pnp, lm_pnp) tasks, suggesting that the autoregressive VLM used in GVL may be biased toward folding-like robotic manipulations. In contrast, VITA achieves consistent performance across all task types and distribution shifts, indicating stronger generalization than in-context learning.
@@ 4.4 Differentiating Expert from Non-Expert Trajectories
Beyond generalization, we evaluate VITA’s robustness by testing its ability to distinguish expert from suboptimal visual trajectories, assigning lower progress scores to the latter. To this end, we compare model predictions on expert and scripted (non-expert) visual trajectories collected from the same in-distribution setting of BridgeData V2. The scripted trajectories are generated in the ToyKitchen environment using a heavily randomized controller [walke2023bridgedata], under the same embodiment and environmental configuration as the expert demonstrations. These include pick-and-place tasks involving object categories that may overlap with the training set, such as general objects (sc_pnp_obj), rigid items (sc_pnp_robj), soft toys (sc_pnp_stoy), utensils (sc_pnp_uten), and a cloth (sc_pnp_cloth). While these trajectories match the training distribution in task, embodiment, and environment, their behavior deviates from the smooth and monotonic progress typically exhibited by expert demonstrations. Appendix C includes examples from the scripted dataset. As scripted datasets are generated by a randomized controller, they are not designed to reflect quantifiable levels of suboptimality. Nonetheless, an effective value function estimator should assign lower VOC scores to suboptimal trajectories relative to expert ones when evaluated in the same setting. We measure discriminative success with BinVOC, a binary evaluation metric defined as $\mathbf{1}\left[\mathrm{VOC}_{\text{exp}}>\mathrm{VOC}_{\text{subopt}}\right]$, which is $1$ when expert trajectories achieve higher VOC than suboptimal ones and 0 otherwise.
@@ 4.5 Zero-Shot Reward Shaping for Offline RL
We evaluate our value function estimator on the Meta-World MT10 benchmark [yu2020meta], which consists of ten diverse robotic manipulation tasks, including pick-and-place, door opening, and pushing. This experiment is designed to evaluate VITA’s zero-shot reward shaping for offline RL. VITA is trained on real-world robotic data and evaluated zero-shot in this simulated setting. For each task in Meta-World MT10 benchmark, we generate 20 expert visual demonstrations using Meta-World’s expert policies. The goal-conditioned value estimator is then used to define a dense reward for each visual observation. Policies are trained with Implicit Q-Learning (IQL) [kostrikov2021offline] in the offline RL setting. We repeat expert data generation and policy training across 10 random seeds. After training, each multi-task policy is evaluated on 20 rollouts per task, and we compute the average return across episodes for each (task, seed) pair. We report the interquartile mean (IQM) across all pairs, together with 95% stratified bootstrap confidence intervals, following the evaluation protocol proposed by agarwal2021deep. The training details and evaluation protocol are further described in Appendix E.
@@ 4.6.1 Effect of Dissimilarity-Based Sampling on Discriminative Performance
We empirically validate the effectiveness of the proposed dissimilarity-based sampling by comparing the impact of different sub-trajectory sampling strategies on VITA’s ability to distinguish expert from non-expert demonstrations. In particular, full-trajectory sampling [sun2024rnn] computes the adaptation loss over all overlapping sub-trajectories with stride $s=1$, which in our setup leads to overfitting to global temporal shortcuts. Random sampling improves diversity by selecting sub-trajectories uniformly, but lacks semantic diversity. In contrast, dissimilarity-based sampling explicitly promotes semantic diversity, resulting in better generalization. Our proposed method, which selects sub-trajectories that maximize pairwise dissimilarity, outperforms both full-trajectory and random sampling in differentiating expert and non-expert visual trajectories. For further details, please refer to Appendix F.1.
@@ 4.6.2 Effect of Temporal Memory on Value Function Estimation
We analyze how different test-time adaptation strategies for incorporating temporal context impact value function estimation. VITA performs step-by-step adaptation without reset (implicit memory), incrementally updating parameters from $\theta_{t-1}$. This allows the value function estimator to encode trajectory history directly into its parameters, capturing both temporal order and accumulated temporal context. We compare VITA with three alternatives: (i) trajectory-level adaptation (TTT-TR), which applies a single gradient update computed over the full trajectory. This produces a batched update that averages the loss across all frames, thereby discarding temporal order; (ii) memoryless adaptation (TTT-RS), which resets the adaptation module to $\theta_{0}$ at every step. By adapting only to the current visual observation, this approach prevents any temporal information from being carried forward; (iii) explicit memory (TTT-EX), which also resets the adaptation module at every step but updates using a local window of recent frames matching the window size used during meta-training ($r=8$). Across all datasets, VITA consistently outperforms trajectory-level, memoryless, and explicit-memory methods, demonstrating the effectiveness of incrementally accumulating temporal context for value function estimation. We provide the full analysis in Appendix F.2.
@@ 5.1 VLMs as Value Function Estimators
Contrastive VLMs have been applied as zero-shot reward models [baumli2023vlm, rocamonde2024vlm] and goal-conditioned value functions [ma2023liv, ma2024vlm], using frame-level similarity scores but lacking temporal modeling. Some works fine-tune CLIP on domain-specific video datasets [fan2022minedojo, jiang2024reinforcement] to generate dense rewards, requiring supervision that limits their zero-shot applicability. RoboCLIP [sontakke2023roboclip] enables one-shot reward specification without finetuning, but yields sparse rewards and still requires demonstrations. Large-scale multimodal pretraining approaches [alakuijala2024video, li2024decisionnce, ma2023liv] learn language-conditioned value functions by relying on scale for generalization. ReWiND [zhang2025rewind] trains a cross-modal sequential transformer on augmented large-scale data with video rewinding, combined with a few in-domain demonstrations to improve generalization. In contrast, VITA adapts a pretrained contrastive vision–language encoder at inference, capturing both temporal and semantic context through test-time adaptation, without requiring large-scale multimodal pretraining or additional domain-specific demonstrations for generalization. VITA could, in principle, be applied to any vision–language representation learning method for decision-making. Autoregressive VLMs [alayrac2022flamingo] leverage in-context learning over trajectories and have been applied to success detection [du2023success] and goal-conditioned value functions [ma2024vlm]. GVL [ma2024vlm] mitigates monotonic bias from ordered pretraining data by shuffling frames at inference, but this discards temporal order, which is essential for temporal reasoning and for distinguishing visually similar yet temporally distinct states. Our approach instead preserves chronological order while capturing temporal context through test-time adaptation, avoiding such shortcuts.
@@ 5.2 Test-Time Adaptation
Test-time adaptation methods provide parameter-efficient solutions for adapting VLMs without the need for full fine-tuning. Prompt-based methods optimize model parameters at test time based on the task context, although their underlying mechanisms are opaque [zhou2022coop, khattak2023maple, ma2023swapprompt]. lim2025sparse suggested that CLIP representations could support test-time adaptation, based on the observation that CLIP encodes human-interpretable concepts discovered via mechanistic interpretability [bricken2023monosemanticity]. CLIP-based similarity has also been used as a reward signal, with parameters optimized at test time via reinforcement learning [zhao2024rlcf]. Another approach is test-time training [sun2020ttt], where an adaptation module is updated at each timestep via a self-supervised loss. Unlike meta-reinforcement learning methods [finn2017maml, bauer2023humantimescale], which require task-specific adaptation episodes during training, test-time training enables direct online adaptation during inference, without the need for task labels.
@@ 6.2 Limitations and Future Work
Test-time adaptation improves generalization of zero-shot progress estimation across unseen tasks and environments in our experiments, but it can still face challenges in settings with high execution variability or extended durations. Although the adaptation overhead is negligible due to the lightweight adaptation module, updating a value function estimator at every timestep may be potentially unsafe during deployment, limiting applicability in real-time scenarios. In future work, we plan to explore alternative approaches for test-time adaptation of vision–language models, particularly in the context of training agents within world models. In addition, applying VITA to real-time closed-loop control settings and complex RL environments remains a promising direction for future work. While we provide empirical evidence that dissimilarity-based sampling mitigates shortcut learning, a theoretical analysis of how diversity-based sampling influences shortcut learning remains a promising direction. This lies beyond the scope of our core contribution, which introduces a value function estimator that meta-learns a test-time adaptation mechanism to improve zero-shot performance and temporal reasoning.
@@ Acknowledgments
This work was supported by UKRI (EP/Y037111/1) as part of the ProSafe project (EU Horizon 2020, MSCA, grant no. 101119358).
@@ D.1 Model and Training Details
We use the OpenCLIP ViT-B/32 encoder pre-trained on the OpenAI dataset as a frozen backbone. Each visual observation and task description is encoded using the CLIP vision and text encoders to produce their representation, respectively. We opted for a joint (concatenation-based) representation over an element-wise product based on improved validation performance. The adaptation module $f_{\text{adapt}}$ is a two-layer residual MLP with GELU activation and a projection dimension $d^{\prime}=64$. The model is trained using the AdamW optimizer with a learning rate of $1\times 10^{-4}$, weight decay of $1\times 10^{-4}$, and a cosine learning rate schedule with 10% warmup. We use a batch size of 32 and pad all trajectories to a maximum length of 120 frames (matching the longest sequence in the training set). The weighting coefficient $\lambda_{\text{self}}$ of the self-supervised loss in the total training objective is set to 0.5, selected based on validation performance. We train for 5 epochs, as validation VOC typically plateaus early, with extended training providing no further improvement. For dissimilarity-based sampling, we set the stride to $s=2$, the sub-trajectory length to $w_{\text{tr}}=8$, and the number of selected windows to $k=16$, unless otherwise specified. All experiments were run on NVIDIA RTX 6000 Ada Generation GPUs.

### v5

@@ 4.2 Baselines
We compare VITA against value functions based on contrastive and autoregressive VLMs. VLM-CL computes zero-shot value estimates from cosine similarity between CLIP representations of each frame and task description (Baumli et al., 2023). VLM-RM regularizes CLIP embeddings by projecting features along the direction from a generic reference prompt to the task prompt (Rocamonde et al., 2024). CLIP-FT trains a supervised regression head on frozen CLIP multimodal representations. CLIP-GRU trains a Gated Recurrent Unit (GRU) (Cho et al., 2014) on frozen CLIP multimodal representations. GVL (Ma et al., 2024) is the state-of-the-art zero-shot value function that leverages autoregressive VLMs. In our experiments, GVL-0S refers to the zero-shot setting, where the VLM is prompted with only the test trajectory and task description. GVL-1S corresponds to a one-shot in-context setting, where a full shuffled trajectory from the training distribution, along with its progress labels, is provided as an example. We follow GVL in using Gemini 1.5 Pro (gemini-1.5 Pro-latest(Gemini Team et al., 2024) with its proposed prompt template. We also tested Qwen-VL 2.5 (Yang et al., 2024), which failed to overcome temporal bias, and GPT-4o (Hurst et al., 2024), which refused to produce value estimates. Additional implementation details are provided in Appendix D.
@@ 4.3 Evaluating Generalization Under Distribution Shifts
We evaluate the ability of progress estimators to generalize across novel environments, tasks, and robot embodiments using subsets of BridgeData V2 curated to introduce variation along these axes. Environment shifts involve changes in scene layout or background. For example, lm_pnp is a pick-and-place task in front of a laundry machine, while td_fold, ft_fold, and rd_fold feature cloth folding on different surfaces. ms_sweep introduces a long-horizon sweeping task in a confined tray. Embodiment shifts are evaluated using the DeepThought robot, which differs from the training embodiment (WidowX 250) in both morphology and camera perspective. dt_tk_pnp (pick-and-place) and dt_tk_stack (stacking) retain the in-distribution environment with a new embodiment, while dt_ft_stack (stacking) and dt_rd_pnp (drawer pick-and-place) involve both embodiment and environment shifts. Each dataset includes 200 expert trajectories with task descriptions, except for ms_sweep, which contains 100 due to its size limitation. Appendix B provides a full description of our evaluation datasets. We evaluate the performance of a value function estimator using the Value Order Correlation (VOC) metric (Ma et al., 2024), which measures the alignment between the predicted progress towards task completion and the chronological order of the frames in a visual trajectory. Formally, let $p_{1},\dots,p_{T}$ denote the predicted progress values for a trajectory of $T$ frames, and let $r_{t}=t$ represent the temporal index. VOC is defined as the Spearman rank correlation $\rho_{s}$ between the predicted values and the temporal indices.
@@
In Table 1, the results show that both VITA and CLIP-GRU consistently outperform GVL-0S and GVL-1S, highlighting the limitations of pre-trained autoregressive VLMs and the importance of preserving temporal order for effective progress estimation. While both demonstrate effective progress estimation, VITA outperforms CLIP-GRU on the majority of evaluation datasets (6 out of 10). Both VITA and CLIP-GRU outperform CLIP-based baselines that lack temporal modeling (CLIP-FT, VLM-CL, VLM-RM), indicating the impact of encoding temporal history on progress estimation. CLIP-FT performs comparably to GVL on the in-distribution task but fails to generalize to out-of-distribution settings. Both VLM-CL and VLM-RM perform poorly, likely due to their lack of temporal modeling. For GVL-0S and GVL-1S, in some cases, in-context examples improve performance, while in others, they provide no benefit or even degrade performance. Both GVL methods perform well on folding tasks (td_fold, ft_fold, rd_fold), but fail to achieve comparable performance on stacking (dt_tk_stack, dt_ft_stack) and pick-and-place (dt_tk_pnp, lm_pnp) tasks, suggesting that the autoregressive VLM used in GVL may be biased toward folding-like robotic manipulations. In contrast, VITA achieves consistent performance across all task types and distribution shifts, indicating stronger generalization than in-context learning.
@@ 4.4 Differentiating Expert from Non-Expert Trajectories
Beyond generalization, we evaluate VITA’s robustness by testing its ability to distinguish expert from suboptimal visual trajectories, assigning lower progress scores to the latter. To this end, we compare model predictions on expert and scripted (non-expert) visual trajectories collected from the same in-distribution setting of BridgeData V2. The scripted trajectories are generated in the ToyKitchen environment using a heavily randomized controller (Walke et al., 2023), under the same embodiment and environmental configuration as the expert demonstrations. These include pick-and-place tasks involving object categories that may overlap with the training set, such as general objects (sc_pnp_obj), rigid items (sc_pnp_robj), soft toys (sc_pnp_stoy), utensils (sc_pnp_uten), and a cloth (sc_pnp_cloth). While these trajectories match the training distribution in task, embodiment, and environment, their behavior deviates from the smooth and monotonic progress typically exhibited by expert demonstrations. Appendix C includes examples from the scripted dataset. As scripted datasets are generated by a randomized controller, they are not designed to reflect quantifiable levels of suboptimality. Nonetheless, an effective value function estimator should assign lower VOC scores to suboptimal trajectories relative to expert ones when evaluated in the same setting. We measure discriminative success with BinVOC, a binary evaluation metric defined as $\mathbf{1}\left[\mathrm{VOC}_{\text{exp}}>\mathrm{VOC}_{\text{subopt}}\right]$, which is $1$ when expert trajectories achieve higher VOC than suboptimal ones and 0 otherwise.
@@ 4.5 Zero-Shot Reward Shaping for Offline RL
We evaluate our value function estimator on the Meta-World MT10 benchmark (Yu et al., 2020), which consists of ten diverse robotic manipulation tasks, including pick-and-place, door opening, and pushing. This experiment is designed to evaluate VITA’s zero-shot reward shaping for offline RL. VITA is trained on real-world robotic data and evaluated zero-shot in this simulated setting. For each task in Meta-World MT10 benchmark, we generate 20 expert visual demonstrations using Meta-World’s expert policies. The goal-conditioned value estimator is then used to define a dense reward for each visual observation. Policies are trained with Implicit Q-Learning (IQL) (Kostrikov et al., 2021) in the offline RL setting. We repeat expert data generation and policy training across 10 random seeds. After training, each multi-task policy is evaluated on 20 rollouts per task, and we compute the average return across episodes for each (task, seed) pair. We report the interquartile mean (IQM) across all pairs, together with 95% stratified bootstrap confidence intervals, following the evaluation protocol proposed by Agarwal et al. (2021). The training details and evaluation protocol are further described in Appendix E.
@@ 4.6.1 Effect of Dissimilarity-Based Sampling on Discriminative Performance
We empirically validate the effectiveness of the proposed dissimilarity-based sampling by comparing the impact of different sub-trajectory sampling strategies on VITA’s ability to distinguish expert from non-expert demonstrations. In particular, full-trajectory sampling (Sun et al., 2024) computes the adaptation loss over all overlapping sub-trajectories with stride $s=1$, which in our setup leads to overfitting to global temporal shortcuts. Random sampling improves diversity by selecting sub-trajectories uniformly, but lacks semantic diversity. In contrast, dissimilarity-based sampling explicitly promotes semantic diversity, resulting in better generalization. Our proposed method, which selects sub-trajectories that maximize pairwise dissimilarity, outperforms both full-trajectory and random sampling in differentiating expert and non-expert visual trajectories. For further details, please refer to Appendix F.1.
@@ 4.6.2 Effect of Temporal Memory on Value Function Estimation
We analyze how different test-time adaptation strategies for incorporating temporal context impact value function estimation. VITA performs step-by-step adaptation without reset (implicit memory), incrementally updating parameters from $\theta_{t-1}$. This allows the value function estimator to encode trajectory history directly into its parameters, capturing both temporal order and accumulated temporal context. We compare VITA with three alternatives: (i) trajectory-level adaptation (TTT-TR), which applies a single gradient update computed over the full trajectory. This produces a batched update that averages the loss across all frames, thereby discarding temporal order; (ii) memoryless adaptation (TTT-RS), which resets the adaptation module to $\theta_{0}$ at every step. By adapting only to the current visual observation, this approach prevents any temporal information from being carried forward; (iii) explicit memory (TTT-EX), which also resets the adaptation module at every step but updates using a local window of recent frames matching the window size used during meta-training ($w_{\text{tr}}=8$). Across all datasets, VITA consistently outperforms trajectory-level, memoryless, and explicit-memory methods, demonstrating the effectiveness of incrementally accumulating temporal context for value function estimation. We provide the full analysis in Appendix F.2.
@@ 5.1 VLMs as Value Function Estimators
Contrastive VLMs have been applied as zero-shot reward models (Baumli et al., 2023; Rocamonde et al., 2024) and goal-conditioned value functions (Ma et al., 2023b; 2024), using frame-level similarity scores but lacking temporal modeling. Some works fine-tune CLIP on domain-specific video datasets (Fan et al., 2022; Jiang et al., 2024) to generate dense rewards, requiring supervision that limits their zero-shot applicability. RoboCLIP (Sontakke et al., 2023) enables one-shot reward specification without fine-tuning, but yields sparse rewards and still requires demonstrations. Large-scale multimodal pre-training approaches (Alakuijala et al., 2024; Li et al., 2024; Ma et al., 2023b) learn language-conditioned value functions by relying on scale for generalization. ReWiND (Zhang et al., 2025b) trains a cross-modal sequential transformer on augmented large-scale data with video rewinding, combined with a few in-domain demonstrations to improve generalization. In contrast, VITA adapts a pre-trained contrastive vision–language encoder at inference, capturing both temporal and semantic context through test-time adaptation, without requiring large-scale multimodal pre-training or additional domain-specific demonstrations for generalization. VITA could be applied to any pre-trained vision–language representation that encodes a universal value function to enhance its generalization and temporal reasoning via meta-learned test-time adaptation. Autoregressive VLMs (Alayrac et al., 2022) leverage in-context learning over trajectories and have been applied to success detection (Du et al., 2023) and goal-conditioned value functions (Ma et al., 2024). GVL (Ma et al., 2024) mitigates monotonic bias from ordered pre-training data by shuffling frames at inference, but this discards temporal order, which is essential for temporal reasoning and for distinguishing visually similar yet temporally distinct states. Our approach instead preserves chronological order while capturing temporal context through test-time adaptation, avoiding such shortcuts.
@@ 5.2 Test-Time Adaptation
Test-time adaptation methods provide parameter-efficient solutions for adapting VLMs without the need for full fine-tuning. Prompt-based methods optimize model parameters at test time based on the task context, although their underlying mechanisms are opaque (Zhou et al., 2022; Khattak et al., 2023; Ma et al., 2023a). Lim et al. (2025) suggested that CLIP representations could support test-time adaptation, based on the observation that CLIP encodes human-interpretable concepts discovered via mechanistic interpretability (Bricken et al., 2023). CLIP-based similarity has also been used as a reward signal, with parameters optimized at test time via reinforcement learning (Zhao et al., 2024). Another approach is test-time training (Sun et al., 2020), where an adaptation module is updated at each timestep via a self-supervised loss. Unlike meta-reinforcement learning methods (Finn et al., 2017; Bauer et al., 2023), which require task-specific adaptation episodes during training, test-time training enables direct online adaptation during inference, without the need for task labels.
@@ 6.2 Limitations and Future Work
Test-time adaptation improves generalization of zero-shot progress estimation across unseen tasks and environments in our experiments, but it can still face challenges in settings with high execution variability or extended durations. Although the adaptation overhead is negligible due to the lightweight adaptation module, updating a value function estimator at every timestep may be potentially unsafe during deployment, limiting applicability in real-time scenarios. In future work, we plan to explore alternative approaches for test-time adaptation of vision–language models, particularly in the context of training agents within world models. In addition, applying VITA to real-time closed-loop control settings and complex RL environments remains a promising direction for future work. VITA could enhance generalization and temporal reasoning for any domain-specific pre-trained multimodal representation via meta-learned test-time adaptation, as it is agnostic to the underlying vision–language encoder. While we provide empirical evidence that dissimilarity-based sampling mitigates shortcut learning, a theoretical analysis of how diversity-based sampling influences shortcut learning remains a promising direction. This lies beyond the scope of our core contribution, which introduces a value function estimator that meta-learns a test-time adaptation mechanism to improve zero-shot performance and temporal reasoning.
@@ Acknowledgments
We thank Daniel Furelos-Blanco, Roko Parac, and Frederik Kelbel for their helpful comments on an earlier draft. We also thank Vassilis Digalakis for valuable discussions. This work was supported by UKRI (EP/Y037111/1) as part of the ProSafe project (EU Horizon 2020, MSCA, grant no. 101119358).
@@ D.1 Model and Training Details
We use the OpenCLIP ViT-B/32 encoder pre-trained on the OpenAI dataset as a frozen backbone. Each visual observation and task description is encoded using the CLIP vision and text encoders to produce their representation, respectively. We opted for a joint (concatenation-based) representation over an element-wise product based on improved validation performance. The adaptation module $f_{\text{adapt}}$ is a two-layer residual MLP with GELU activation and a projection dimension $d^{\prime}=64$. The model is trained using the AdamW optimizer with a learning rate of $1\times 10^{-4}$, weight decay of $1\times 10^{-4}$, and a cosine learning rate schedule with 10% warmup. We use a batch size of 32 and pad all trajectories to a maximum length of 120 frames (matching the longest sequence in the training set). The weighting coefficient $\lambda_{\text{self}}$ of the self-supervised loss in the total training objective is set to 0.5, selected based on validation performance. We train for 5 epochs, as validation VOC typically plateaus early, with extended training providing no further improvement. For dissimilarity-based sampling, we set the stride to $s=1$, the sub-trajectory length to $w_{\text{tr}}=8$, and the number of selected windows to $k=8$, unless otherwise specified. All experiments were run on NVIDIA RTX 6000 Ada Generation GPUs.